OpenStudy (anonymous):

Calc2 evaluating the definite integral of x sqrt (1+2x) from 0 to 4, keep getting around 7.5, but should be around 20 integrating by parts, u=x, dv=sqrt (1+2x)dx, v=2/3(1=2x)^(3/2), du=dx end up with 2/3 (x)(1+2x)^(3\2) - 4/15 (1+2x)^(5/2) so good so far or did I miss something?

OpenStudy (anonymous):

why not just a simple u-substitution? \(\large u=1+2x\rightarrow du=2dx\rightarrow \frac{1}{2}du=dx \) to replace the remaining x, use \(\large \frac{(u-1)}{2}=x \)

OpenStudy (anonymous):

I would love to, but instructions say to integrate by parts...

OpenStudy (anonymous):

lol... you have to use IBP ?

OpenStudy (anonymous):

yes, just to make it difficult

OpenStudy (anonymous):

do you also have to write it out using your non-writing hand?

OpenStudy (anonymous):

just kidding....

OpenStudy (anonymous):

no thank goodness

OpenStudy (anonymous):

ok... integration by parts....

OpenStudy (anonymous):

i see it... your dv=(1+2x)^(1/2) so v = (1/3)(1+2x)^(3/2)

OpenStudy (anonymous):

\[2/3x(1+2x)^{3/2}-\int\limits_{0}^{4}2/3(1=2X)^{3/2}dx\]

OpenStudy (anonymous):

why is dv=(1+2x)^3/2dx?

OpenStudy (anonymous):

sorry... that should be a 1/2...

OpenStudy (anonymous):

\[\huge dv=(1+2x)^{\frac{1}{2}}dx\rightarrow v=\frac{1}{3}(1+2x)^{\frac{3}{2}} \]

OpenStudy (anonymous):

shouldn't that be 2/3(bla bla)? 1/3/2?

OpenStudy (anonymous):

same as 2/3?

OpenStudy (anonymous):

no because when you take the derivative, of v, you get (1+2x)^(1/2)... i think you may be missing the chain rule...

OpenStudy (anonymous):

I think your right, I don't like chains

OpenStudy (anonymous):

so your integral should be.,..

OpenStudy (anonymous):

\[\large u\cdot v - \int v du = x\cdot \frac{1}{3}(1+2x)^{\frac{3}{2}}-\int \frac{1}{3}(1+2x)^{\frac{3}{2}}dx \]

OpenStudy (anonymous):

would the integral of \[x \sqrt{x-1}\] be much different?

OpenStudy (anonymous):

as far as v?

OpenStudy (anonymous):

not much different....

OpenStudy (anonymous):

in fact, that would be easier since you don't need to think about the chain rule in this case...

OpenStudy (anonymous):

book shows that as \[2/3 (x-1)^{3/2}\] so i am a little confused

OpenStudy (anonymous):

is that v or dv?

OpenStudy (anonymous):

v

OpenStudy (anonymous):

that's correct because if you integrate \(\large dv=\sqrt{x-1}dx \), \(\large v=\int \sqrt{x-1}dx=\frac{2}{3}(x-1)^{\frac{3}{2}}\)

OpenStudy (anonymous):

Ok, so I have to look at the chain rule some more...

OpenStudy (anonymous):

yeah... technically this second one does the chain rule but the derivative of the "inside" is 1....

OpenStudy (anonymous):

chain says f(g(x))=f'(g(x))*g'(x) for derivatives,

OpenStudy (anonymous):

outside then inside right?

OpenStudy (anonymous):

yep... do you understand what is meant by derivative of the "inside" function?

OpenStudy (anonymous):

inside is what is in the ( )... in this case the derivative of the inside is 2?

OpenStudy (anonymous):

(1+2x)

OpenStudy (anonymous):

yes... the derivative of the "inside" function is 2 in that original question posted...

OpenStudy (anonymous):

that's it^^^

OpenStudy (anonymous):

ok so it is 2/3(bla bla)*2, and I was missing the *2 i think, no cuz thats 4/3

OpenStudy (anonymous):

Hang on let me think about this for a sec

OpenStudy (anonymous):

1 mississippi

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

2 mississippi

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

\[\int\limits_{?}^{?}(1+2x)^{1/2}dx=2/3(1+2x)^{3/2}*2\]? close?

OpenStudy (anonymous):

still 4/3

OpenStudy (anonymous):